OpenStudy (anonymous):

The probability of winning a particular game at the Game Place is 1/4 . Find the odds against winning the game.

OpenStudy (anonymous):

P(Winning the game) = 1/4 P(Losing the game) = 1 - 1/4 = ?

OpenStudy (anonymous):

3/4?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

thankyou very much (: